OpenStudy (anonymous):

Simplify 4^3 sqrt 125

OpenStudy (anonymous):

\[4^{3}\sqrt{125} = 4^{3}\sqrt{(5^{2}\times5)} = 4^{3}\times5\sqrt{5}\]